OpenStudy (anonymous):

A gas is expanded from an initial volume of 20.5 L at 0.92 atm at 23 °C to a final volume of 34.6 L. During the expansion the gas cools to 12 °C. What is the final pressure

OpenStudy (anonymous):

Use the equation: \[\frac{ PV }{ T } = \frac{ PV }{ T }\]

OpenStudy (anonymous):

Be sure to convert °C to Kelvin (K) when writing T in the equation.

OpenStudy (anonymous):

Ok so when substituting the values with the formula you provided i got this: 0.92 atm* 20.5L/ 296 K = P * 34.6L/ 285 K This much i understand, how do i solve for the pressure in the second equation?

OpenStudy (anonymous):

Calculate the number and unit on the left, calculate P * the number and unit on the right. To solve for P, use your algebraic skills. Divide by the number on the right to isolate the variable P. Your Liters and Kelvin cancel and you're left with an answer with atm as a pressure unit.